OpenStudy (anonymous):

What is the value of x? 4(x+2)=14-2(3-4x)

OpenStudy (anonymous):

the first step is to distribute.

OpenStudy (anonymous):

4x+8=14-6-8x

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now you gotta combine like terms

OpenStudy (anonymous):

4x+8=8-8x

OpenStudy (anonymous):

good. now you gotta get the Xs only on one side

OpenStudy (anonymous):

12x+8=8

OpenStudy (anonymous):

now you have to cancel out the 8

OpenStudy (anonymous):

12x=16?

OpenStudy (anonymous):

wrong. you have to subtract 8 from one side

OpenStudy (anonymous):

12x=0

OpenStudy (anonymous):

The answer matter what your teacher wants. You can answer it as undefined or no solution

OpenStudy (anonymous):

so it would just be zero????

OpenStudy (anonymous):

no the answer would not be a number. It would be no solution or undefined.

OpenStudy (anonymous):

it has to be a number.... thats all i can put into the computer is numbers

OpenStudy (anonymous):

huh ....... :K Bump up the question to see if anyone else knows.

OpenStudy (anonymous):

x=0, but 4x+8=14-6+8x; 4x-8x=14-6-8; -4x=0; x=0